OpenStudy (anonymous):

Write an equation of a line in slope intercept form going through the points (2, 1) and (- 1, - 5). Show your work

OpenStudy (anonymous):

first we need to find the slope(m)... m = (y2 - y1) / (x2 - x1) x1 = 2, x2 = -1 y1 = 1, y2 = -5 lets sub it in... m = (-5 - 1) / (-1 - 2) m = -6/-3 = 2 so our slope is 2 y - y1 = m(x - x1) using slope 2 and points (2,1) y - 1 = 2(x - 2) y - 1 = 2x - 4 y = 2x - 4 + 1 y = 2x - 3 <-- your equation

OpenStudy (anonymous):

i love your brain lol ^~^

OpenStudy (anonymous):

I love my brain too ...lol :)